OpenStudy (anonymous):

A supervisor has determined that the average salary of the employees in his department is $40000 with a standard deviation of $15000. A sample of 25 of the employees salaries was selected at random. Assuming the distribution of the salaries is normal, what is the probability that the average for this sample is between $36000 and $42000?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

I have no clue how to solve this problem.

OpenStudy (anonymous):

\[\bar{X}~N(40000,(\frac{15000^2}{25})\]\[\frac{15000^2}{25}=9\times10^6\]\[P(\bar{X}<42000)=P(Z<\frac{42000-40000}{\sqrt{9\times10^6}})\approx P(Z<0.67)=\Phi(0.67)=0.7486\]Therefore :\[P(\mu<\bar{X}<42000)\approx0.7486-0.5=0.2486\]\[P(36000<\bar{X}<\mu)=P(\bar{X}<44000)-0.5=P(Z<\frac{44000-400000}{\sqrt{9\times10^6}})-0.5\]\[\approx P(Z<1.33)-0.5\approx \Phi(1.33)-0.5=0.9082-0.5=0.4082\]So:\[P(36000<\bar{X}<42000)\approx 0.2486+0.4082=0.6568\]

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

God, that took a long time! :P

OpenStudy (anonymous):

haha ty i have been staring at that problem for 2 days.......dont go anywhere i got another one for ya lol. im aaron btw

OpenStudy (anonymous):

I hope it's nothing that takes as long as an estimation question, I'm tired now :P

OpenStudy (anonymous):

lol i know right

OpenStudy (anonymous):

By the way, I'm sorry if a mistake in that :P I'm guessing you haven't seen one if you haven't said.

OpenStudy (anonymous):

A tire manufacturer claims that the lives of its tires are normally distributed and will last an average of 60,000 miles with a standard deviation of 3000 miles. 64 randomly selected tires were tested and the average miles where the tire failed were recorded. What is the probability that the mean miles recorded when the tires failed will be more than 59,500 miles?

OpenStudy (anonymous):

it was correct lol

OpenStudy (anonymous):

How do you know? You have the answers?

OpenStudy (anonymous):

Yes it tells me if I have it worked out correctly.

OpenStudy (anonymous):

the new one looks just as confusing lol

OpenStudy (anonymous):

Arghh, it just deleted all my work but basically, I get 0.9082.

OpenStudy (anonymous):

lol i hate when that happens

OpenStudy (anonymous):

Do you live in America?

OpenStudy (anonymous):

that was correct, lol ok one more problem you have just literally saved me 10 hours im terrible at this..........yes i live in st. louis mo

OpenStudy (anonymous):

Ok, I'm assuming your in ~12th grade, yeah? I strongly recommend this book: http://www.amazon.com/Statistics-AQA-SMP-AS-Mathematics/dp/052160527X/ref=sr_1_1?ie=UTF8&s=books&qid=1307222760&sr=8-1

OpenStudy (anonymous):

The average life of a particular battery is determined to be 56.2 hours. With a standard deviation of 4.3 hours. Assume battery life is normally distributed. Batteries in the lower 10% do not meet minimum life requirements. What is the minimum life requirement for a battery?

OpenStudy (anonymous):

Damn! I accidentally pressed refresh and it deleted all my work again. 50.69?

OpenStudy (anonymous):

yeppers, can you help me on a few more? i would appreciate it

OpenStudy (anonymous):

Ok, I could do with practising these anyway :)

OpenStudy (anonymous):

Ages of applicants for a particular scholarship are normally distributed with a mean age of 21.3 years and a standard deviation of 1.85 years. If an applicant is randomly selected, find the probability that the applicant is under 18 years old.

OpenStudy (anonymous):

still here?

OpenStudy (anonymous):

\[X~N(21.3,1.85^2)\implies P(X<18)=1-P(X<24.6)=1-P(Z<\frac{24.6-21.3}{1.85})\]\[\approx1-P(Z<1.78)=1-\Phi(1.78)=1-0.9625=0.0375\]

OpenStudy (anonymous):

sweet rock on dont go anywhere lol

OpenStudy (anonymous):

so the final answer was .04?

OpenStudy (anonymous):

Yeah, but normally, in statistics, we give more significant figures than that.

OpenStudy (anonymous):

ok thats cool were supposed to round up to 2 decimals

OpenStudy (anonymous):

Oh, that's somewhat ridiculous but whatever.

OpenStudy (anonymous):

lol yeah i know

OpenStudy (anonymous):

Song lengths scheduled for airplay on a radio station ar normally distributed with a mean length of 3.5 minutes and a standard deviation of .28 minute. What percentage of songs is within 3 to 4 minutes long?

OpenStudy (anonymous):

This is making so much more sense I appreciate it.

OpenStudy (anonymous):

That makes it worthwhile.

OpenStudy (anonymous):

\[X~(3.5,0.28)\implies P(3<X<\mu)=P(\mu<X<4)=P(X<4)-0.5\] \[=P(Z<\frac{4-3.5}{0.28})-0.5\approx P(Z<1.79)-0.5=\Phi(1.79)-0.4\]\[=0.9633-0.5=0.4633\]\[\implies P(3<X<4)=2\times0.4633=0.9266\]

OpenStudy (anonymous):

Can I go to bed now?

OpenStudy (anonymous):

one more please lol

OpenStudy (anonymous):

Go on then.

OpenStudy (anonymous):

3.) The average life of a particular bacteria strain is determined to be 27.3 hours with a standard deviation of 2.6 hours. Assume bacteria life is normally distributted. Bacteria in the lower 10% are determined to be inconsequential.l What is the minimum life required for studay?

OpenStudy (anonymous):

23.97 hrs

OpenStudy (anonymous):

i actually have one more if you dont mind staying realy quick

OpenStudy (anonymous):

i over looked it

OpenStudy (anonymous):

Ok, lets see it.

OpenStudy (anonymous):

k just a sec :)

OpenStudy (anonymous):

A catering company claims that the total calories of their entrees are normally distributed with an average of 1000 calories and a standard deviation of 300 calories. Fifty randomly selected meals were tested and the average calories were recorded. What is the probability that the mean calories recorded is greater than 1100 calories?

OpenStudy (anonymous):

there you go :)

OpenStudy (anonymous):

\[X~N(1000,300^2)\implies \bar{X}~N(1000, \frac{300^2}{50})\Longleftrightarrow \bar{X}~N(1000, 1800)\]\[\implies P(\bar{X}>1100)=1-P(\bar{X}<1100)=1-P(Z<\frac{1100-1000}{\sqrt{1800}})\]\[\approx P(Z<2.36)=\Phi(Z)=0.9909\]

OpenStudy (anonymous):

hey are you usually on here? thanks for your help!

OpenStudy (anonymous):

Damn forgot to do 1-ans, it should be 1-0.9909=0.0091

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Yeah, I'm normally on here, feel free to ask me any time :)

OpenStudy (anonymous):

ok great thank you! :)

OpenStudy (anonymous):

That's fine.

OpenStudy (anonymous):

hello are you here today?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Hi.

OpenStudy (anonymous):

Hello are you here?

OpenStudy (anonymous):

Just post a question and I'll answer it when I'm next online :)

OpenStudy (anonymous):

1,) A university surveys 36 randomly selected students and asks the number of dollars spent on textbooks/supplies during the academic year. The results are (in dollars are shown below. 425 78 53 699 663 282 845 373 161 22 853 653 366 806 920 562 828 931 641 134 916 38 379 90 388 894 654 639 916 284 700 684 600 425 717 822 Construct a 90% confidence interval for the mean dollars spent. Would a given mean value of $600 be likely, given the calculated confidence level?

OpenStudy (anonymous):

aaron.vallero@yahoo.com is my email

OpenStudy (anonymous):

Hi are you here?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

For that confidence interval, can you work out the mean for me? And do you know what the unbiased estimator S is?

OpenStudy (anonymous):

i believe the mean is the $600 but thats all i understand with it

OpenStudy (anonymous):

Do you know how to calculate the mean?

OpenStudy (anonymous):

I thought it was the sum of all the numbers

OpenStudy (anonymous):

The mean is the sum of all the numbers divided by the amount of numbers there are. S is what we use to estimate the standard deviation of a population from a sample, it is given by:\[S=\frac{\sum_{i=1}^{n}x _{i}^2-n \bar{x}^2}{n-1}\]Are you ok with this notation?

OpenStudy (anonymous):

Yeah thats fine

OpenStudy (anonymous):

Ok, so now we've estimated the standard deviation, \(\bar{X}\) follows a t distribution with n-1 (which in this case is 35) degrees of freedom. If we then look up 95% in the percentage points of the t distribution with 35 degrees of freedom, this will give us the t value we need to use, which is 1.69 (use 95% because you want 5% each side). So the interval is:\[\bar{x}\pm 1.69\frac{\S}{\sqrt{n}}\approx540\pm 1.69\frac{294.96}{\sqrt{36}}\approx540\pm83.1\]Using the proper notation:\[(456.9,623.1)\]

OpenStudy (anonymous):

$600 is in the interval and therefore, seems possible. Oh, I made one mistake, what I said was \(S\) was actually \(S^2\) so you need to take the square root to get \(S\). Please say if any of this is unclear.

OpenStudy (anonymous):

hey are you here?

OpenStudy (anonymous):

??